OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. Should be two answers. |4x + 3| = 9 + 2x

Parth (parthkohli):

\[|4x + 3| = 9 + 2x\implies 4x + 3 = \pm(9 +2x) \]

Parth (parthkohli):

So, we have two equations that we solve individually.\[ 4x + 3 = 9 + 2x\]\[4x + 3 = -9 - 2x\]

OpenStudy (anonymous):

i don't know..lol

Parth (parthkohli):

lol...\[4x + 3 = 9 + 2x \]Subtract 3 from both sides.\[4x = 6 + 2x\]Subtract 2x from both.\[2x = 6\]Divide both sides by 2.\[x = {6 \over 2}=3\]

Parth (parthkohli):

Similarly, solve the second equation.